A cohort study of physical activity and coronary heart disease was conducted in a rural community. The incidence of heart disease was 18.1 per 1,000 person-years among those with sedentary lifestyles and 9.1 per 1,000 person-years among those with active lifestyles. We know that 35% of this population lived sedentary lifestyles. What is the proportion of risk in the total population attributable to a sedentary lifestyle?
Answer #1

Proportion of risk attributable to Sedentary lifestyle:

Attributable proportion also known as Attributable risk % or Attributable fraction . The formula for calculating Attributable proportion or Attributable risk percentage is:

AP= RD/ Cle = Cle- Clu/ Cle

Attributable proportion or Attributable percentage can be calculated by taking the difference ( Risk difference, RD) between the incidence in the risk group and non risk group ( sedentary group and active group,Cle - Clu) and then divide the answer with the incidence of risk group. Risk percentage can be calculated by multiplying the answer with 100.

Attributable Proportion = RD/ Cle = (18.1- 9.1)/ 18.1 = 9/ 18.1= 0.497

Attributablerisk percentage =0.497× 100= 49.7 %

It can be concluded as 49.7% of heart disease in the total population is attributed to Sedentary lifestyle.